# Discovery of a nova super-remnant surrounding the recurrent nova LMCN 1971-08a in the Large Magellanic Cloud

### v1

@@
A nova super-remnant (NSR) is a greatly-extended structure grown by repeated nova eruptions sweeping the surrounding material away from a nova into a dense outer shell and are predicted to form around all novae. To date, four NSRs are known, with three located in the Galaxy and one residing in M31. Here we present the discovery of the first NSR in the Large Magellanic Cloud and only the second extragalactic nova shell to be identified, hosted by the recurrent nova LMCN 1971-08a. The structure is coincident with the nova, has a circular morphology and is visible in narrowband H and filters but very faint in , as expected. Hi data also potentially reveal the existence of a coincident structure. Further, with a diameter of pc, this NSR is the largest example yet found, with models indicating an shell expanding at into the surrounding medium and an age of 2.4 Myr. The existence of the NSR also suggests that LMCN 1971-08a may have a much shorter recurrence period than currently presumed.
@@
Recurrent novae (RNe) are observationally-defined as accreting white dwarf (WD) binaries that have exhibited more than one nova eruption. Such eruptions are driven by a thermonuclear runaway on the WD’s surface (Starrfield et al., 1972) that leads to a portion of the accreted material being ejected at velocities up to thousands of kilometers per second (O’Brien et al., 2001) that, in some cases, form expanding sub-parsec nova shells (Shara et al., 2024b; Santamaría et al., 2025). After the eruption, the resumption of accretion onto the WD (Worters et al., 2007) prepares the nova system for subsequent eruptions. Furthermore, accumulation of material retained on the WD following each eruption can grow the WD (Yaron et al., 2005; Hillman et al., 2016) to the Chandrasekhar limit to potentially explode as a type Ia supernova (Whelan & Iben, 1973).
@@
Nova super-remnants (NSRs) are shell-like structures orders of magnitude larger than singular eruption nova shells (for a review see Healy-Kalesh, 2024). The first NSR (Darnley et al., 2019) was discovered in the Andromeda Galaxy around the rapidly recurring nova, M31N 2008-12a, and has been formed by the ejecta of repeated nova eruptions over millenia excavating tens of thousands of solar masses of surrounding interstellar medium from around the central binary into a high density shell (Darnley et al., 2019). Theoretically, such structures should encompass all novae (Healy-Kalesh et al., 2023) though their size on the sky and low surface brightnesses make NSRs difficult to find in optical data. However, with the capabilities of the recently deployed Condor Array Telescope (Lanzetta et al., 2023), optical NSRs have now been discovered around the Galactic RNe KT Eridani (Shara et al., 2024a; Healy-Kalesh et al., 2024c), T Coronae Borealis (Shara et al., 2024c) and RS Ophiuchi (Shara et al., 2025)111A possible cavity component of the NSR associated with RS Ophiuchi was identified in far-infrared data (Healy-Kalesh et al., 2024a).. These NSRs have diameters ranging from pc; have emission most prominent in H, and ; lack strong emission as a result of ionisation from low velocity shocks; and exhibit low expansion velocities in their outer shells. While the NSR around M31N 2008-12a appears to have a more well-defined elliptical shell as it is viewed pole on (Darnley et al., 2019), the Galactic NSRs have more patchy morphologies likely as a result of the NSR growing within the irregular interstellar medium (ISM) and possibly from being observed at different viewing angles. Furthermore, the NSRs surrounding T CrB (Shara et al., 2024c) and RS Oph (Shara et al., 2025) appear bilobed, possibly resulting from the ejecta being initially influenced by the accretion disk around the WD. A search for more NSRs in M31 and the Large Magellanic Cloud (LMC) did not lead to any more examples being found (Healy-Kalesh et al., 2024b) but this apparent dearth was attributed to the structures around other ‘younger’ systems being considerably fainter.
@@
LMCN 1971-08a is one of the four currently known recurrent novae in the LMC. It exhibited eruptions in 1971 (Graham, 1971) and 2009 (Liller, 2009) and therefore has a recurrence period (or inter-eruption interval) of /N years (Bode et al., 2016), where N may be one or a small integer. While only a spectrum of the first eruption was obtained in 1971 with no further follow-up (Graham, 1971; Bode et al., 2016), the more recent eruption in 2009 was followed in great detail, being observed in the optical, UV, near-infrared (Bode et al., 2016) and X-rays (Bode et al., 2016; Orio et al., 2021). These observations revealed a very-fast declining nova comprising a WD accreting material at an elevated rate of from a sub-giant companion with a probable orbital period of 1.2 days (Bode et al., 2016). The ejecta velocities seen from the nova ranged from (Orio et al., 2009; Bode et al., 2016) with shocks likely present in the ejecta before day nine of the eruption (Bode et al., 2016).
@@
In this Letter, we present the discovery of a nova super-remnant around the recurrent nova LMCN 1971-08a; the first nova super-remnant to be found in the LMC. In Section 2 we outline the datasets used in this work before describing the characteristics of the NSR across optical and radio data in Section 3. In Section 4 we place the discovery within context of the growing field of nova super-remnants and rule out other possible origins before summarising our findings in Section 5.
@@ 2 Optical and radio data
Within this study, several datasets were utilised across the optical and radio. The Magellanic Cloud Emission Line Survey (MCELS; Smith & Team, 1999) observed the central of the LMC with five filters including narrow-band H, and filters. The DeMCELS survey (Points et al., 2024) is a contemporary higher-resolution emission-line survey of the Magellanic Clouds undertaken with the Dark Energy Camera (DECam; Flaugher et al., 2015) narrow-band H and filters. The continuum-subtracted narrowband data from both MCELS and DeMCELS were utilised for this work. The HI4PI survey (HI4PI Collaboration et al., 2016) is an all-sky neutral atomic hydrogen (Hi) survey constructed from the Effelsberg-Bonn Hi Survey (EBHIS; Kerp et al., 2011) and the Galactic All-Sky Survey (GASS; McClure-Griffiths et al., 2009) data sets, thereby providing better angular resolution and sensitivity than previous all-sky Hi surveys. The MeerKAT radio data (Cotton et al. in prep.) used in this work are from the MeerKAT array (Jonas & MeerKAT Team, 2016), a new generation radio telescope with high resolution and sensitivity such that it is capable of resolving structures in the LMC (see, e.g. Sasaki et al., 2025).
@@ 3 LMCN 1971-08a nova super-remnant
A coherent, circular shell-like structure spatially coincident with LMCN 1971-08a appears in the MCELS H data (see Figure 1), with the same shell visible in the MCELS image but negligible in MCELS , as is expected for NSRs (Darnley et al., 2019; Shara et al., 2024a, c, 2025). While clearly evident in the MCELS data, the high resolution DeMCELS H and imaging of the NSR emphasizes the structure further, as shown in Figure 2 (with the approximate extent of the NSR indicated).
With a distance of kpc to the LMC (Pietrzyński et al., 2019), the NSR has a projected diameter of approximately 200 pc. As seen in the optical data of Figure 1, the NSR is brighter to the northeast and southwest with a fainter boundary connecting the two components to the northwest, defining the NSR outer shell edge. The shell thickness is taken to be from the southwest edge as the inner and outer boundaries of this part of the NSR shell are pc and pc from the nova, respectively. This possible compression front may indicate, as with a similar feature in the KT Eri NSR, interaction of a bow shock with an earlier phase of the remnant’s evolution (Healy-Kalesh et al., 2024c). Additionally, there appears to be an apparent interaction with a complex structure to the west (Figure 3) composed of many Hii regions (Pellegrini et al., 2012).
Following Points et al. (2024), the H and surface brightnesses of the bright northeast component of the NSR are and , respectively. The H and surface brightnesses of the bright southwest component are marginally fainter radiating at and , respectively.
@@ 3.1 NSR in radio data
As shown in Figure 1, within the Hi data cube from the HI4PI survey (HI4PI Collaboration et al., 2016) sliced at and , there does appear to be correlated structure coincident with the H shell. Specifically, there appears to be a higher level of emission from the southwestern portion of the NSR in the slice and a possible shell boundary to the northeast in the slice. A velocity distribution across the NSR with a range of in the declination and right ascension direction (see Figure 4 for the slits used) does reveal an arc with a range of as shown in Figure 5, possibly indicating an outer shell expansion velocity of . Within MeerKAT radio data (Cotton et al. in prep.), there does not strongly appear to be a corresponding structure, as shown in Figure 1. By estimating the RMS noise level in the region of the NSR shell, we derive an upper limit of Jy / beam.
@@ 3.2 Simple model of NSR evolution
To estimate the surrounding ISM density for the model, we took the Hi data velocity cube from the HI4PI survey (HI4PI Collaboration et al., 2016), converted the integrated intensity in the range of into column density (using equation 2 from HI4PI Collaboration et al., 2016) and found the average within of the nova (to incorporate the NSR and its surroundings). Assuming a depth of 5 kpc (estimated from the NSR’s location in Figure 9 of Subramanian & Subramaniam, 2009) gives an ISM density of (number density of ).
With a nova system mimicking LMCN 1971-08a in a prepopulated ISM estimated for the local region, it would take years over 62,400 eruptions to reach a radius of 100 pc, containing an evacuated cavity region 15 pc in radius. The NSR outer shell would have a mass of (made up almost exclusively of ISM) and an expansion velocity of , which is consistent with the expansion velocity derived from the Hi position-velocity analysis from Section 3.1.
@@ 4.1 Comparison to other NSRs
With a diameter of pc, the LMCN 1971-08a NSR is larger than the NSR around M31N 2008-12a (major axis of 134 pc; Darnley et al., 2019) and a factor of , and larger than the NSRs associated with KT Eri (50 pc; Shara et al., 2024a), T CrB (30 pc; Shara et al., 2024c) and RS Oph (70 pc; Shara et al., 2025), respectively. The projected morphology of the NSR is also more circular than the other NSRs, likely resulting from the system’s inclination, and is most alike the M31N 2008-12a NSR with more defined edges than the irregular morphology of the Galactic NSRs.
As with all other NSRs, this NSR is most prominent through its H and emission. The H surface brightnesses of the brightest part of the NSR (northeast component) is . Having a H surface brightness comparable to the other NSRs: for KT Eri (Shara et al., 2024a), for T CrB (Shara et al., 2024c), for RS Oph (Shara et al., 2025) and for M31N 2008-12a (based on integrated H flux; Darnley et al. 2019), this NSR is intrinsically brighter than all other NSRs, supporting predictions that larger NSRs should be brighter (Healy-Kalesh et al., 2023). Moreover, the NSR shell is almost negligible through the narrowband filter (Figure 7) due to the lack of photoionisation across the structure, as found for the other NSRs (Darnley et al., 2019; Shara et al., 2024a, c, 2025), though the southwest component does have very faint emission coinciding with a higher /H ratio.
The Hi position-velocity analysis indicates a possible expansion velocity of the NSR outer edges to be . These velocities are consistent with those expansion velocities seen in other NSRs shells (Darnley et al., 2019; Shara et al., 2024a, 2025) and fits with the range of NSR outer shell velocities predicted from modelling in this work and other studies (e.g. Healy-Kalesh et al., 2023).
Both LMCN 1971-08a and another NSR-hosting nova KT Eri show ”several remarkable similarities” (Bode et al., 2016). Those authors suggested both novae may belong to a class of RNe with WDs approximately away from the Chandrasekhar limit but with very high mass transfer rates between eruptions. This similarity may indicate a potential correlation between these types of novae and the brightest NSRs, but cannot be tested further until the sample of NSRs increases.
@@ 4.2 Other origins
It is necessary to consider other origins of the visible shell encompassing LMCN 1971-08a including the structure being a supernova remnant (SNR), a superbubble, a planetary nebula (PN) or a Hii region. We therefore searched for sources within the NSR ( radius) using SIMBAD (Wenger et al., 2000) and derived emission line ratios (/H, /H and /[S ii]; see Section A.3) for consideration of other possible scenarios.
While the structure is bright in H and , the low levels of emission (see Section A.3 and Figure 7) and lack of radio emission (see MeerKAT panel in Figure 1) from the shell (both typical of SNR shock fronts) combined with zero radio sources found within the structure, low expansion velocities inferred from Hi data (see Section 3.1) and it being larger than any SNR in the LMC (e.g. Zangrandi et al., 2024) demonstrate that the discovered structure is not a SNR. The absence of strong emission (see Figure 7) and the size of the shell (LMC PNe are unresolved in MCELS data as they are arcsec-scale; e.g. Shaw et al., 2001) rules out a PN origin. There is one young stellar object (Whitney et al., 2008), one cluster of stars (Bica et al., 2008) and two associations of stars (Bica et al., 2008) within the shell but their classification does not elaborate as to whether they are OB associations (Bica et al., 2008), which could be responsible for ionising the local ISM and growing a bubble of this size (Wright, 2020). Also, the structure is not included in the Hii region catalogue by Pellegrini et al. (2012) so rules against it being a Hii region or a superbubble, as does the lack of radio emission and relatively high /H ratio (see Figure 8).

### v2

@@
A nova super-remnant (NSR) is a greatly-extended structure grown by repeated nova eruptions sweeping the surrounding material away from a nova into a dense outer shell and are predicted to form around all novae. To date, four NSRs are known, with three located in the Galaxy and one residing in M31. Here we present the discovery of the first NSR in the Large Magellanic Cloud and only the second extragalactic nova shell identified, hosted by the recurrent nova LMCN 1971-08a. The structure is coincident with the nova, has a circular morphology, and is visible in narrowband H and filters but is very faint in , as expected. Hi data also potentially reveal the existence of a coincident structure. Further, with a diameter of pc, this NSR is the largest example yet found, with models indicating an shell expanding at into the surrounding medium and an age of 2.4 Myr. The existence of the NSR also suggests that LMCN 1971-08a may have a much shorter recurrence period than currently presumed.
@@
Recurrent novae (RNe) are observationally-defined as accreting white dwarf (WD) binaries that have exhibited more than one nova eruption. Such eruptions are driven by a thermonuclear runaway on the WD’s surface (Starrfield et al., 1972), which leads to a portion of the accreted material being ejected at velocities up to thousands of kilometers per second (O’Brien et al., 2001); in some cases, this material forms expanding sub-parsec nova shells (Shara et al., 2024b; Santamaría et al., 2025). After the eruption, the resumption of accretion onto the WD (Worters et al., 2007) prepares the nova system for subsequent eruptions. Furthermore, accumulation of material retained on the WD following each eruption can grow the WD (Yaron et al., 2005; Hillman et al., 2016) to the Chandrasekhar limit to potentially explode as a type Ia supernova (Whelan & Iben, 1973).
@@
Nova super-remnants (NSRs) are shell-like structures orders of magnitude larger than singular eruption nova shells (for a review see Healy-Kalesh, 2024). The first NSR (Darnley et al., 2019) was discovered in the Andromeda Galaxy around the rapidly recurring nova, M31N 2008-12a and formed by the ejecta of repeated nova eruptions over millennia, excavating tens of thousands of solar masses of surrounding interstellar medium from around the central binary into a high-density shell (Darnley et al., 2019). Theoretically, such structures should encompass all novae (Healy-Kalesh et al., 2023), although their size on the sky and low surface brightnesses make NSRs difficult to find in optical data. However, with the capabilities of the recently deployed Condor Array Telescope (Lanzetta et al., 2023), optical NSRs have now been discovered around the Galactic RNe KT Eridani (Shara et al., 2024a; Healy-Kalesh et al., 2024c), T Coronae Borealis (Shara et al., 2024c) and RS Ophiuchi (Shara et al., 2025)111A possible cavity component of the NSR associated with RS Ophiuchi was identified in far-infrared data (Healy-Kalesh et al., 2024a).. These NSRs have diameters ranging from pc; have emission most prominent in H, and ; lack strong emission as a result of ionisation from low velocity shocks; and exhibit low expansion velocities in their outer shells. While the NSR around M31N 2008-12a appears to have a more well-defined elliptical shell as it is viewed pole-on (Darnley et al., 2019), the Galactic NSRs have more patchy morphologies likely as a result of the NSR growing within the irregular interstellar medium (ISM) and possibly from being observed at different viewing angles. Furthermore, the NSRs surrounding T CrB (Shara et al., 2024c) and RS Oph (Shara et al., 2025) appear bilobed, possibly resulting from the ejecta being initially influenced by the accretion disk around the WD. A search for additional NSRs in M31 and the Large Magellanic Cloud (LMC) did not yield further examples (Healy-Kalesh et al., 2024b), but this apparent dearth was attributed to the structures around other ”younger” systems being considerably fainter.
@@
LMCN 1971-08a is one of the four currently known recurrent novae in the LMC. It exhibited eruptions in 1971 (Graham, 1971) and 2009 (Liller, 2009) and therefore has a recurrence period (or inter-eruption interval) of /N years (Bode et al., 2016), where N may be one or a small integer. While only a spectrum of the first eruption was obtained in 1971 with no further follow-up (Graham, 1971; Bode et al., 2016), the more recent eruption in 2009 was followed in great detail, being observed in the optical, UV, near-infrared (Bode et al., 2016), and X-rays (Bode et al., 2016; Orio et al., 2021). These observations revealed a very fast-declining nova comprising a WD, accreting material at an elevated rate of from a subgiant companion with a probable orbital period of 1.2 days (Bode et al., 2016). The ejecta velocities seen from the nova ranged from (Orio et al., 2009; Bode et al., 2016) with shocks likely present in the ejecta before day nine of the eruption (Bode et al., 2016).
@@
In this Letter, we present the discovery of a nova super-remnant around the recurrent nova LMCN 1971-08a, the first nova super-remnant to be found in the LMC. In Sect. 2 we outline the datasets used in this work before describing the characteristics of the NSR across optical and radio data in Sect. 3. In Sect. 4 we place the discovery within the context of the growing field of nova super-remnants and rule out other possible origins before summarising our findings in Sect. 5.
@@ 2 Optical and radio data
Within this study, several datasets were utilised across the optical and radio. The Magellanic Cloud Emission Line Survey (MCELS; Smith & Team, 1999) observed the central of the LMC with five filters including narrowband H, and filters. The DeMCELS survey (Points et al., 2024) is a contemporary higher-resolution emission-line survey of the Magellanic Clouds undertaken with the Dark Energy Camera (DECam; Flaugher et al., 2015) narrowband H and filters. The continuum-subtracted narrowband data from both MCELS and DeMCELS were utilised for this work. The HI4PI survey (HI4PI Collaboration et al., 2016) is an all-sky neutral atomic hydrogen (Hi) survey constructed from the Effelsberg-Bonn Hi Survey (EBHIS; Kerp et al., 2011) and the Galactic All-Sky Survey (GASS; McClure-Griffiths et al., 2009) data sets, which provide better angular resolution and sensitivity than previous all-sky Hi surveys. The MeerKAT radio data (Cotton et al., in preparation) used in this work are from the MeerKAT array (Jonas & MeerKAT Team, 2016), a new generation radio telescope with high resolution and sensitivity such that it is capable of resolving structures in the LMC (see, e.g., Sasaki et al., 2025).
@@ 3 LMCN 1971-08a nova super-remnant
A coherent, circular shell-like structure spatially coincident with LMCN 1971-08a appears in the MCELS H data (see Fig. 1), with the same shell visible in the MCELS image but negligible in MCELS , as expected for NSRs (Darnley et al., 2019; Shara et al., 2024a, c, 2025). While clearly evident in the MCELS data, the high resolution DeMCELS H and imaging of the NSR emphasizes the structure further, as shown in Fig. 2 (with the approximate extent of the NSR indicated).
With a distance of kpc to the LMC (Pietrzyński et al., 2019), the NSR has a projected diameter of 200 pc. As seen in the optical data of Fig. 1, the NSR is brighter to the northeast and southwest, with a fainter boundary connecting these two components to the northwest, defining the outer shell edge of the NSR. The shell thickness is taken to be from the southwest edge, as the inner and outer boundaries of this part of the NSR shell are pc and pc from the nova, respectively. This possible compression front may indicate interaction of a bow shock with an earlier phase of the remnant’s evolution, as seen with a similar feature in the KT Eri NSR, (Healy-Kalesh et al., 2024c). Additionally, there appears to be an interaction with a complex structure to the west (Fig. 3) composed of many Hii regions (Pellegrini et al., 2012).
Following Points et al. (2024), the H and surface brightnesses of the bright northeast component of the NSR are and , respectively. The H and surface brightnesses of the bright southwest component are marginally fainter, radiating at and , respectively.
@@ 3.1 NSR in radio data
As shown in Fig. 1, the Hi data cube from the HI4PI survey (HI4PI Collaboration et al., 2016) sliced at and , reveals a correlated structure coincident with the H shell. Specifically, the southwestern portion of the NSR shows a higher level of emission from in the slice, while a possible shell boundary appears to the northeast in the slice. A velocity distribution across the NSR, spanning in declination and right ascension (see Fig. 4 for the slits used) reveals an arc with a range of , as shown in Fig. 5, possibly indicating an outer shell expansion velocity of . In MeerKAT radio data (Cotton et al., in preparation), there does not appear to be a corresponding structure, as shown in Fig. 1. By estimating the RMS noise level in the region of the NSR shell, we derive an upper limit of Jy / beam.
@@ 3.2 Simple model of NSR evolution
To estimate the surrounding ISM density for the model, we took the Hi data velocity cube from the HI4PI survey (HI4PI Collaboration et al., 2016), converted the integrated intensity in the range of into column density (using Eq. 2 from HI4PI Collaboration et al., 2016), and found the average within of the nova (to incorporate the NSR and its surroundings). Assuming a depth of 5 kpc (estimated from the NSR’s location in Fig. 9 of Subramanian & Subramaniam, 2009) gives an ISM density of (number density of ).
With a nova system mimicking LMCN 1971-08a in a pre-populated ISM estimated for the local region, it would take years over 62,400 eruptions to reach a radius of 100 pc, containing an evacuated cavity region 15 pc in radius. The NSR outer shell would have a mass of (made up almost exclusively of ISM) and an expansion velocity of , which is consistent with the expansion velocity derived from the Hi position-velocity analysis from Sect. 3.1.
@@ 4.1 Comparison to other NSRs
With a diameter of pc, the LMCN 1971-08a NSR is 1.5 times larger than the NSR around M31N 2008-12a (major axis of 134 pc; Darnley et al., 2019) and a factor of 4, 3, and 6.5 times larger than the NSRs associated with KT Eri (50 pc; Shara et al., 2024a), T CrB (30 pc; Shara et al., 2024c) and RS Oph (70 pc; Shara et al., 2025), respectively. The projected morphology of the NSR is also more circular than the other NSRs, likely resulting from the system’s inclination, and is most similar to the M31N 2008-12a NSR with more defined edges than the irregular morphology of the Galactic NSRs.
As with all other NSRs, this NSR is most prominent through its H and emission. The H surface brightnesses of the brightest component of the NSR (northeast part) is . Having an H surface brightness comparable to the other NSRs – for KT Eri (Shara et al., 2024a), for T CrB (Shara et al., 2024c), for RS Oph (Shara et al., 2025), and for M31N 2008-12a (based on integrated H flux; Darnley et al. 2019) – this NSR is intrinsically brighter than all others, supporting predictions that larger NSRs should be brighter (Healy-Kalesh et al., 2023). Moreover, the NSR shell is almost negligible in the narrowband filter (Fig. 7) due to the lack of photoionisation across the structure, consistent with the findings of other NSRs (Darnley et al., 2019; Shara et al., 2024a, c, 2025). However, the southwest part does exhibit very faint emission coinciding with a higher /H ratio.
The Hi position-velocity analysis indicates a possible expansion velocity of the NSR outer edges of . These velocities are consistent with those expansion velocities seen in other NSRs shells (Darnley et al., 2019; Shara et al., 2024a, 2025) and are in agreement with the range of NSR outer shell velocities predicted by modelling in this work and other studies (e.g., Healy-Kalesh et al., 2023).
Both LMCN 1971-08a and another NSR-hosting nova KT Eri exhibit ”several remarkable similarities” (Bode et al., 2016). Those authors suggested that both novae may belong to a class of RNe with WDs approximately below the Chandrasekhar limit, but with very high mass transfer rates between eruptions. This similarity may indicate a potential correlation between these types of novae and the brightest NSRs, although this cannot be tested further until the sample of known NSRs increases.
@@ 4.2 Other origins
It is necessary to consider alternative origins of the visible shell encompassing LMCN 1971-08a, including it being a supernova remnant (SNR), a superbubble, a planetary nebula (PN), or an Hii region. We therefore searched for sources within the NSR ( radius) using SIMBAD (Wenger et al., 2000) and derived emission line ratios (/H, /H, and /[S ii]; see Appendix A.3) for a consideration of these alternative scenarios.
While the structure is bright in H and , the low levels of emission (see Appendix A.3 and Fig. 7) and lack of radio emission (see MeerKAT panel in Fig. 1) from the shell (both typical of SNR shock fronts), the absence of radio sources within the structure, the low expansion velocities inferred from Hi data (see Sect. 3.1), and its size exceeding that of any SNR in the LMC (e.g., Zangrandi et al., 2024) together demonstrate that the discovered structure is not an SNR. The absence of strong emission (see Fig. 7) and the size of the shell (LMC PNe are unresolved in MCELS data as they are arcsec-scale; e.g., Shaw et al., 2001) rule out a PN origin. There is one young stellar object (Whitney et al., 2008), one cluster of stars (Bica et al., 2008), and two associations of stars (Bica et al., 2008) within the shell. However, their classification does not specify whether they are OB associations (Bica et al., 2008), which could ionise the local ISM and create a bubble of this size (Wright, 2020). Also, the structure is absent from the Hii region catalogue by Pellegrini et al. (2012), arguing against it being an Hii region or a superbubble. This is further supported by the lack of radio emission and the relatively high /H ratio (see Fig. 8).